I know when factoring the formula is ax^2(+-)bx(+-)c is the formula I should use. But, what do I do when I'm given an equation such as this? 6x^3-4x^2-2x or x^4-10x^2+9. I have been working on this equation for the past hour-------4X^4-38x^3+48x^2. None of these equations follow the formula, and yet they should. Can anyone explain it to me?
Answer by josgarithmetic(39617)

You hope to factor: 6x^3-4x^2-2x
Look for common factors among the terms. See, 2x. What you have after reverse of distributive property may still be factorable, too.



You also hope to factor: x^4-10x^2+9
That is certainly factorable according to what you were trying to do with quadratic trinomials. Do what you already know. You see, some number like, t=x^2 ? You see 9=9*1, 10=-9+(-1) ?